OpenStudy (j2lie):

How do you solve 8x^2+20x-48?

OpenStudy (whpalmer4):

By "solve" do you mean find the values of \(x\) such that \[8x^2+20x-48 = 0\]

OpenStudy (whpalmer4):

Or are you trying to factor the polynomial?

OpenStudy (anonymous):

To find the x's, you have to factor.

OpenStudy (whpalmer4):

No, you don't have to factor, but that is one way you can find them.

OpenStudy (whpalmer4):

In fact, you can find the values of \(x\) even if the polynomial is irreducible (cannot be factored).

OpenStudy (j2lie):

You have to factor it. How do you do that?

OpenStudy (whpalmer4):

Is your assignment to factor it, or find the values of \(x\) that make it equal to 0?

OpenStudy (whpalmer4):

@Thefaceless uh, that's not correct...

OpenStudy (whpalmer4):

\[8x^2+20x - 48 = 2*2*2*x^2+2*2*5x-2*2*2*2*3\]so you can factor out 2*2 = 4 giving you\[4(2x^2+5x-12)\]

OpenStudy (anonymous):

Oh whoops forgot + and 2

OpenStudy (whpalmer4):

@j2lie one last time: are you supposed to factor that polynomial, or solve it for the values of \(x\) that make it equal to 0?

OpenStudy (anonymous):

She said factor

OpenStudy (anonymous):

use the quadratic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] @j2lie

OpenStudy (whpalmer4):

Okay, if you are factoring it, we already got out the common factor of 4: \[4(2x^2+5x-12)\] Now we are going to have two binomials, one of which will be \((x+a)\) for some value of \(a\) to be determined, and the other of which will be \((2x + b\), for some value of \(b\) to be determined. We know that one of them must have a \(2\) because we need to end up with \(2x^2\) as our highest order term. Let's try an experiment: \[(x+a)(2x+b) = 2x^2 + bx + 2ax + ab\]\[=2x^2 + (2a+b)x + ab\] Compare with our polynomial to be factored:\[2x^2+5x-12\] To make those the same, we need to find values of \(a,b\) such that\[2a+b=5\]\[a*b=-12\] Can you think of two numbers that satisfy those conditions?

OpenStudy (whpalmer4):

Obviously, one will be positive and the other negative. I'll list the factors of 12: 1*12 2*6 3*4 4*3 6*2 12*1 If you put a - sign in front of one number in the pair, can you find the pair that works?

OpenStudy (anonymous):

quadratic formula is much faster :3

OpenStudy (j2lie):

i figured it out. but I got stuck.

OpenStudy (whpalmer4):

Yes, it is, but it is important to know how to factor, and that's what she says she is supposed to do in this problem.

OpenStudy (anonymous):

oh. if she doesn't know that, guess she's ought to do it the hard way.

OpenStudy (j2lie):

you're supposed to split the middle but it doesn't work.

OpenStudy (whpalmer4):

how did you split the middle?

OpenStudy (j2lie):

-3x and 8x

OpenStudy (anonymous):

try & use this once if you still can't figure it out: \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (whpalmer4):

ah, but you have to divide one by 2 because of the multiplier...

OpenStudy (whpalmer4):

@sayakdbz countless millions of algebra students have learned how to factor trinomials. j2lie can do it too :-)

OpenStudy (anonymous):

@whpalmer4 i'll leave this one to u :)

OpenStudy (j2lie):

i am confused

OpenStudy (whpalmer4):

Okay. do you agree that the factoring will look something like \[(2x + a)(x+b)\]?

OpenStudy (j2lie):

yes

OpenStudy (whpalmer4):

Good. So if you "split the middle", don't you have to divide one of the pieces in half? It's going to be multiplied by 2, not 1.

OpenStudy (anonymous):

u are guiding it in the wrong way. its supposed to be: \[x^2+(a+b)x+ab=(x+a)*(x+b)\] @whpalmer4

OpenStudy (whpalmer4):

@sayakdbz how many polynomials do you think I've factored in the 35 years I've been doing this?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

@thefaceless 2, actually, this will be the 3rd :-)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

XD

OpenStudy (whpalmer4):

if you're factoring by grouping: \[ax^2+bx+c\]you multiply a*c, and choose two factors of the product a*c that sum to b. -8 and 3 are such a pair. \[2x^2 -8x + 3x - 12\]Group terms:\[(2x^2-8x) + (3x-12)\]Now factor each group. what do you get?

OpenStudy (whpalmer4):

\[(2x^2-8x)+(3x-12) = 2x(x-4) + 3(x-4)\]Notice anything interesting about the right hand side?

OpenStudy (whpalmer4):

Sorry, we're not doing \(2x^2-5x+12\) we're doing \(2x^2+5x-12\)! -3 and 8 are our split of the middle: \[2x^2 + 8x -3x -12\]\[(2x^2+8x)-(3x+12)\]\[2x(x+4)-3(x+4)\] What do we get when we factor \((x+4)\) out of both of those terms? Essentially using the distributive property in reverse... \[2x(x+4)+3(x+4) = (x+4)(2x+3)\] Combine that with the common factor of \(4\) that we pulled out earlier: \[8x^2+20x-48 = 4(x+4)(2x-3)\] If we want to "solve" that (find the values of \(x\) that make the value of that be \(0\), we can set \[4(x+4)(2x-3) = 0\]\[(x+4)=0,\,x=-4\]\[2x-3=0,\,2x=3,\,x=3/2\] So \(x=-4\) and \(x = 3/2\) are the solutions.

OpenStudy (j2lie):

the solution is (x+4)(8x-12)

OpenStudy (whpalmer4):

What do you think \(4(2x-3)\) = ?

OpenStudy (j2lie):

8x-12

OpenStudy (whpalmer4):

As I said several lines earlier: \[8x^2+20x-48 = 4(x+4)(2x-3)\]That is the complete factoring. \[(x+4)(8x-12)\]is an incomplete factoring, because there is still a factor of \(4\) which can be taken out of the \((8x-12)\) product term. "Solving" a polynomial is commonly understood to mean finding the roots, aka zeros, aka values that make the polynomial = 0.

OpenStudy (j2lie):

I understand what you mean. but it's from 8x^2+32x_12x-48.

OpenStudy (whpalmer4):

\[8x^2+32x-12x-48 = 8x^2 + 20x - 48\]What's your point? If you're factoring that by grouping: \[(8x^2+32x) - (12x+48) = 8x(x+4)-12(x+4)\]\[ =(8x-12) (x+4)\]\[ = 4(2x-3)(x+4)\]

OpenStudy (j2lie):

I did it this way 8x^2+20x-48. 48*8=384 8x^2+32x-12x-48 8x(x+4)-12(x+4)= 8x^2+20x-48

OpenStudy (whpalmer4):

That's all good except for the last line. You need to factor out the \((x+4)\) from both terms — that's why you did the splitting/grouping.

OpenStudy (j2lie):

whoops I forgot but I did that.

OpenStudy (whpalmer4):

Do you recognize that \(8x(x+4) - 12(x+4)\) is simply a case of \[b*a + c*a\]where \(a = (x+4)\)?

OpenStudy (whpalmer4):

and then we can rewrite it as \[a(b+c)\]

OpenStudy (j2lie):

yes it's like a foiling problem.

OpenStudy (whpalmer4):

so \[8x(x+4) - 12(x+4)\]can be written as \[(x+4)*8x+(x+4)*(-12) = (x+4)(8x-12)\]which can be further factored to \[4(x+4)(2x-3)\]

OpenStudy (j2lie):

to me it's easier leaving it that way because the 4 is confusing.

OpenStudy (whpalmer4):

okay, but it isn't fully factored without pulling out the 4. And in many applications, you'll want it fully factored so you can cancel common factors (say this is in the numerator of a fraction, and the denominator contains \((2x-3)\) — will you realize that you can cancel it out?)

OpenStudy (j2lie):

yes

OpenStudy (whpalmer4):

Anything else you need to know about this problem?

OpenStudy (j2lie):

i did know how to factor the 4 with the terms in the parentheses.

OpenStudy (whpalmer4):

Yes, I'm sure you do — the question is whether you will notice it in the middle of a large problem when it isn't an obvious number to factor out. You want to stack the deck in your favor. Trusting that you'll spot a common factor like the 4 in this case isn't doing that.

OpenStudy (j2lie):

I did it you're way and the answer is the same.

OpenStudy (j2lie):

@whpalmer4 thank you for helping me.

OpenStudy (whpalmer4):

You're welcome! Sorry if I seemed a bit grumpy — my pizza was late :-)

OpenStudy (j2lie):

it's because I was arguing with you, that's why you're grumpy.